# On a counterexample to Mordell’s Pellian equation conjecture: a non-computer based approach

### v1

@@ 2. Preliminaries
$\bullet$ $a_{1}a_{2}=6890530871592$,
$b_{1}a_{2}+a_{1}b_{2}=-d+21554509784529$.
$\bullet$ $6890530871592a_{3}=-315587d+28244816980645$,
$21554509784529a_{3}+6890530871592b_{3}=-972567d+5178868509861$.
$\bullet$ $28244816980645a_{4}=-1272650d+15513835265240$,
$5178868509861a_{4}+28244816980645b_{4}=2697384d+4467486423321$.
$\bullet$ $15513835265240a_{5}=-675500d+35811027000980$,
$4467486423321a_{5}+15513835265240b_{5}=652149d+30537129644761$.
$\bullet$ $35811027000980a_{6}=-1309781d+20714594938939$,
$30537129644761a_{6}+35811027000980b_{6}=6236863d+16497568105021$.
$\bullet$ $20714594938939a_{7}=-441941d+21223637681572$,
$16497568105021a_{7}+20714594938939b_{7}=28697d+18837954382492$.
$\bullet$ $21223637681572a_{8}=-407041d+13338881989779$,
$18837954382492a_{8}+21223637681572b_{8}=4721606d+17458165870498$.
$\bullet$ $13338881989779a_{9}=-230639d+3661427934964$,
$17458165870498a_{9}+13338881989779b_{9}=5070046d+17002573835617$.
$\bullet$ $3661427934964a_{10}=-52048d+15123266299360$,
$17002573835617a_{10}+3661427934964b_{10}=130052d+34143299535672$.
$\bullet$ $15123266299360a_{11}=-202084d+22867483171996$,
$34143299535672a_{11}+15123266299360b_{11}=6208881d+10414810162433$.
$\bullet$ $22867483171996a_{12}=-282499d+8083910734589$,
$10414810162433a_{12}+22867483171996b_{12}=7832662d+17492146626900$.
$\bullet$ $8083910734589a_{13}=-92186d+37797607362521$,
$17492146626900a_{13}+8083910734589b_{13}=1575020d+23141664104777$.
$\bullet$ $37797607362521a_{14}=-411842d+28944089196110$,
$23141664104777a_{14}+37797607362521b_{14}=1939038d+6609389335501$.
$\bullet$ $28944089196110a_{15}=-239446d+16720583049264$,
$6609389335501a_{15}+28944089196110b_{15}=2568152d+34718516820815$.
$\bullet$ $16720583049264a_{16}=-120376d+6912281518232$,
$34718516820815a_{16}+16720583049264b_{16}=1211229d+10227378309170$.
$\bullet$ $6912281518232a_{17}=-44790d+5195414457466$,
$10227378309170a_{17}+6912281518232b_{17}=741048d+289652693936$.
$\bullet$ $5195414457466a_{18}=-30152d+22465854532740$,
$289652693936a_{18}+5195414457466b_{18}=265509d+17460040771835$.
$\bullet$ $22465854532740a_{19}=-123878d+18660863010082$,
$17460040771835a_{19}+22465854532740b_{19}=2001631d+19162392372661$.
$\bullet$ $18660863010082a_{20}=-99441d+31621822715453$,
$19162392372661a_{20}+18660863010082b_{20}=1834171d+26774899887696$.
$\bullet$ $31621822715453a_{21}=-167069d+18945559551357$,
$26774899887696a_{21}+31621822715453b_{21}=4282685d+14809237415832$.
$\bullet$ $18945559551357a_{22}=-93807d+25490759313159$,
$14809237415832a_{22}+18945559551357b_{22}=1035530d+16642252647821$.
$\bullet$ $25490759313159a_{23}=-82380d+31410566401668$,
$16642252647821a_{23}+25490759313159b_{23}=5180714d+7665156542464$.
$\bullet$ $31410566401668a_{24}=-97899d+24750475711101$,
$7665156542464a_{24}+31410566401668b_{24}=4386896d+12345706811136$.
$\bullet$ $24750475711101a_{25}=-34513d+4341133124105$,
$12345706811136a_{25}+24750475711101b_{25}=8807546d+33705248273090$.
$\bullet$ $4341133124105a_{26}=-5385d+20420906899575$,
$33705248273090a_{26}+4341133124105b_{26}=888730d+19138455539665$.
$\bullet$ $20420906899575a_{27}=-21329d+11628413963866$,
$19138455539665a_{27}+20420906899575b_{27}=432256d+14371072675706$.
$\bullet$ $11628413963866a_{28}=-10322d+27536499676132$,
$14371072675706a_{28}+11628413963866b_{28}=2230673d+20483449818325$.
$\bullet$ $27536499676132a_{29}=-16214d+11871313871146$,
$20483449818325a_{29}+27536499676132b_{29}=10491877d+23841147446157$.
$\bullet$ $11871313871146a_{30}=1168d+36708968638606$,
$23841147446157a_{30}+11871313871146b_{30}=583202d+29824187122383$.
$\bullet$ $36708968638606a_{31}=27670d+35292993605984$,
$29824187122383a_{31}+36708968638606b_{31}=6287182d+30828612734725$.
$\bullet$ $35292993605984a_{32}=33791d+32099367907223$,
$30828612734725a_{32}+35292993605984b_{32}=2928174d+15667131309718$.
$\bullet$ $32099367907223a_{33}=31502d+8686497691888$,
$15667131309718a_{33}+32099367907223b_{33}=1722251d+13069916858845$.
$\bullet$ $8686497691888a_{34}=17842d+32178145784978$,
$13069916858845a_{34}+8686497691888b_{34}=1025368d+38021671372563$.
$\bullet$ $32178145784978a_{35}=68542d+38082283375434$,
$38021671372563a_{35}+32178145784978b_{35}=5937837d+11732682959923$.
$\bullet$ $38082283375434a_{36}=100145d+35965182833267$,
$11732682959923a_{36}+38082283375434b_{36}=6202247d+37211608563518$.
$\bullet$ $35965182833267a_{37}=106375d+33169240097520$,
$37211608563518a_{37}+35965182833267b_{37}=8936948d+26020836879171$.
$\bullet$ $33169240097520a_{38}=99259d+17708055977779$,
$26020836879171a_{38}+33169240097520b_{38}=10150666d+32544641385778$.
$\bullet$ $17708055977779a_{39}=70781d+30770548150680$,
$32544641385778a_{39}+17708055977779b_{39}=2202939d+36915064492287$.
$\bullet$ $30770548150680a_{40}=155955d+11904176271315$,
$36915064492287a_{40}+30770548150680b_{40}=2562988d+165258901717$.
$\bullet$ $11904176271315a_{41}=64624d+25508831080099$,
$165258901717a_{41}+11904176271315b_{41}=1347305d+5907415462461$.
$\bullet$ $25508831080099a_{42}=143454d+872140076252$,
$5907415462461a_{42}+25508831080099b_{42}=790654d+34225185690977$.
$\bullet$ $872140076252a_{43}=5137d+10987048845161$,
$34225185690977a_{43}+872140076252b_{43}=339647d+26138731918415$.
$\bullet$ $10987048845161a_{44}=74345d+8152289751913$,
$26138731918415a_{44}+10987048845161b_{44}=1871482d+38377823723539$.
$\bullet$ $8152289751913a_{45}=56296d+17399000858672$,
$38377823723539a_{45}+8152289751913b_{45}=295750d+32817075202061$.
$\bullet$ $17399000858672a_{46}=140874d+13424492118010$,
$32817075202061a_{46}+17399000858672b_{46}=2189576d+33163475056782$.
$\bullet$ $13424492118010a_{47}=116689d+22068166729399$,
$33163475056782a_{47}+13424492118010b_{47}=705436d+24070914532262$.
$\bullet$ $22068166729399a_{48}=266604d+20874064459588$,
$24070914532262a_{48}+22068166729399b_{48}=292240d+15318085049644$.
$\bullet$ $20874064459588a_{49}=267012d+25167892634080$,
$15318085049644a_{49}+20874064459588b_{49}=789145d+30753673516637$.
$\bullet$ $25167892634080a_{50}=384789d+20538358542749$,
$30753673516637a_{50}+25167892634080b_{50}=1701637d+23065592907429$.
$\bullet$ $20538358542749a_{51}=591526d+21236595367747$,
$23065592907429a_{51}+20538358542749b_{51}=3306393d+31259701520158$.
$\bullet$ $21236595367747a_{52}=628169d+663951808753$,
$31259701520158a_{52}+21236595367747b_{52}=1963739d+37377022803680$.
$\bullet$ $663951808753a_{53}=20667d+11352564579816$,
$37377022803680a_{53}+663951808753b_{53}=1565314d+35652369557844$.
$\bullet$ $11352564579816a_{54}=408337d+12677902910401$,
$35652369557844a_{54}+11352564579816b_{54}=6095135d+24252482581547$.
$\bullet$ $12677902910401a_{55}=632652d+37493335194470$,
$24252482581547a_{55}+12677902910401b_{55}=2876153d+28715685383254$.
$\bullet$ $37493335194470a_{56}=1912828d+27624385992608$,
$28715685383254a_{56}+37493335194470b_{56}=11211960d+33470792628624$.
$\bullet$ $27624385992608a_{57}=1505206d+34038147456710$,
$33470792628624a_{57}+27624385992608b_{57}=6487920d$.
@@
$z\equiv$ $\left(0+\sqrt{d}\right)^{3}\left(2+\sqrt{d}\right)^{6}\left(0+\sqrt{d}\right)^{2}\left(1+\sqrt{d}\right)^{3}\left(2+\sqrt{d}\right)^{6}\left(0+\sqrt{d}\right)^{2}\left(0+\sqrt{d}\right)^{0}\left(1+\sqrt{d}\right)^{2}\left(0+\sqrt{d}\right)^{3}$
$\left(2+\sqrt{d}\right)^{6}\left(2+\sqrt{d}\right)^{2}\left(0+\sqrt{d}\right)^{1}\left(2+\sqrt{d}\right)^{5}\left(2+\sqrt{d}\right)^{7}\left(2+\sqrt{d}\right)^{1}\left(1+\sqrt{d}\right)^{3}\left(2+\sqrt{d}\right)^{2}\left(2+\sqrt{d}\right)^{3}$
$\left(0+\sqrt{d}\right)^{1}\left(2+\sqrt{d}\right)^{3}\left(1+\sqrt{d}\right)^{7}\left(0+\sqrt{d}\right)^{3}\left(1+\sqrt{d}\right)^{2}\left(1+\sqrt{d}\right)^{4}\left(1+\sqrt{d}\right)^{0}\left(0+\sqrt{d}\right)^{3}\left(1+\sqrt{d}\right)^{3}$
$\left(0+\sqrt{d}\right)^{3}\left(0+\sqrt{d}\right)^{1}\left(0+\sqrt{d}\right)^{3}\left(1+\sqrt{d}\right)^{7}\left(0+\sqrt{d}\right)^{1}\left(1+\sqrt{d}\right)^{2}\left(1+\sqrt{d}\right)^{3}\left(1+\sqrt{d}\right)^{2}\left(0+\sqrt{d}\right)^{2}$
$\left(1+\sqrt{d}\right)^{7}\left(2+\sqrt{d}\right)^{1}\left(1+\sqrt{d}\right)^{6}\left(2+\sqrt{d}\right)^{5}\left(1+\sqrt{d}\right)^{5}\left(2+\sqrt{d}\right)^{3}\left(2+\sqrt{d}\right)^{3}\left(1+\sqrt{d}\right)^{1}\left(1+\sqrt{d}\right)^{7}$
$\left(2+\sqrt{d}\right)^{0}\left(0+\sqrt{d}\right)^{0}\left(1+\sqrt{d}\right)^{4}\left(1+\sqrt{d}\right)^{2}\left(2+\sqrt{d}\right)^{7}\left(0+\sqrt{d}\right)^{0}\left(2+\sqrt{d}\right)^{7}\left(0+\sqrt{d}\right)^{2}\left(2+\sqrt{d}\right)^{6}$
$\left(2+\sqrt{d}\right)^{7}\left(2+\sqrt{d}\right)^{6}\left(2+\sqrt{d}\right)^{6}$
@@
It follows from Claims 2 and 4 that $\eta\in\mathcal{O}_{K}^{\times}\setminus\{1\}$. Since $\varepsilon^{-d}<\eta<\varepsilon^{d}$, there is some $k\in\mathbb{Z}\setminus\{0\}$ such that $-d<k<d$ and $\eta=\varepsilon^{k}$. Set ${\rm ord}\left(\varepsilon\mathcal{O}_{d}^{\times}\right)=\min\{k\in\mathbb{N}\mid\left(\varepsilon\mathcal{O}_{d}^{\times}\right)^{k}=\mathcal{O}_{d}^{\times}\}$. Observe that ${\rm ord}\left(\varepsilon\mathcal{O}_{d}^{\times}\right)=\left(\mathcal{O}_{K}^{\times}:\mathcal{O}_{d}^{\times}\right)\mid d$ (because $\mathcal{O}_{K}^{\times}/\mathcal{O}_{d}^{\times}$ is generated by $\varepsilon\mathcal{O}_{d}^{\times}$ and $d\in\mathbb{P}$ is ramified in $\mathcal{O}_{K}$). Since $\left(\varepsilon\mathcal{O}_{d}^{\times}\right)^{k}=\eta\mathcal{O}_{d}^{\times}=\mathcal{O}_{d}^{\times}$ by Claim 3, we have that ${\rm ord}\left(\varepsilon\mathcal{O}_{d}^{\times}\right)\mid k$, and thus $\left(\mathcal{O}_{K}^{\times}:\mathcal{O}_{d}^{\times}\right)=1$ (since $d\in\mathbb{P}$). We infer that $\varepsilon\in\mathcal{O}_{d}$, and hence $d\mid y$.∎(Theorem 1.2)

### v2

@@ 2. Preliminaries
$\bullet$ $a_{1}a_{2}=6890530871592$,                                                            $b_{1}a_{2}+a_{1}b_{2}=-d+21554509784529$.
$\bullet$ $6890530871592a_{3}=-315587d+28244816980645$,          $21554509784529a_{3}+6890530871592b_{3}=-972567d+5178868509861$.
$\bullet$ $28244816980645a_{4}=-1272650d+15513835265240$,     $5178868509861a_{4}+28244816980645b_{4}=2697384d+4467486423321$.
$\bullet$ $15513835265240a_{5}=-675500d+35811027000980$,        $4467486423321a_{5}+15513835265240b_{5}=652149d+30537129644761$.
$\bullet$ $35811027000980a_{6}=-1309781d+20714594938939$,     $30537129644761a_{6}+35811027000980b_{6}=6236863d+16497568105021$.
$\bullet$ $20714594938939a_{7}=-441941d+21223637681572$,        $16497568105021a_{7}+20714594938939b_{7}=28697d+18837954382492$.
$\bullet$ $21223637681572a_{8}=-407041d+13338881989779$,        $18837954382492a_{8}+21223637681572b_{8}=4721606d+17458165870498$.
$\bullet$ $13338881989779a_{9}=-230639d+3661427934964$,          $17458165870498a_{9}+13338881989779b_{9}=5070046d+17002573835617$.
$\bullet$ $3661427934964a_{10}=-52048d+15123266299360$,          $17002573835617a_{10}+3661427934964b_{10}=130052d+34143299535672$.
$\bullet$ $15123266299360a_{11}=-202084d+22867483171996$,     $34143299535672a_{11}+15123266299360b_{11}=6208881d+10414810162433$.
$\bullet$ $22867483171996a_{12}=-282499d+8083910734589$,        $10414810162433a_{12}+22867483171996b_{12}=7832662d+17492146626900$.
$\bullet$ $8083910734589a_{13}=-92186d+37797607362521$,          $17492146626900a_{13}+8083910734589b_{13}=1575020d+23141664104777$.
$\bullet$ $37797607362521a_{14}=-411842d+28944089196110$,     $23141664104777a_{14}+37797607362521b_{14}=1939038d+6609389335501$.
$\bullet$ $28944089196110a_{15}=-239446d+16720583049264$,     $6609389335501a_{15}+28944089196110b_{15}=2568152d+34718516820815$.
$\bullet$ $16720583049264a_{16}=-120376d+6912281518232$,        $34718516820815a_{16}+16720583049264b_{16}=1211229d+10227378309170$.
$\bullet$ $6912281518232a_{17}=-44790d+5195414457466$,            $10227378309170a_{17}+6912281518232b_{17}=741048d+289652693936$.
$\bullet$ $5195414457466a_{18}=-30152d+22465854532740$,          $289652693936a_{18}+5195414457466b_{18}=265509d+17460040771835$.
$\bullet$ $22465854532740a_{19}=-123878d+18660863010082$,     $17460040771835a_{19}+22465854532740b_{19}=2001631d+19162392372661$.
$\bullet$ $18660863010082a_{20}=-99441d+31621822715453$,        $19162392372661a_{20}+18660863010082b_{20}=1834171d+26774899887696$.
$\bullet$ $31621822715453a_{21}=-167069d+18945559551357$,     $26774899887696a_{21}+31621822715453b_{21}=4282685d+14809237415832$.
$\bullet$ $18945559551357a_{22}=-93807d+25490759313159$,        $14809237415832a_{22}+18945559551357b_{22}=1035530d+16642252647821$.
$\bullet$ $25490759313159a_{23}=-82380d+31410566401668$,        $16642252647821a_{23}+25490759313159b_{23}=5180714d+7665156542464$.
$\bullet$ $31410566401668a_{24}=-97899d+24750475711101$,        $7665156542464a_{24}+31410566401668b_{24}=4386896d+12345706811136$.
$\bullet$ $24750475711101a_{25}=-34513d+4341133124105$,          $12345706811136a_{25}+24750475711101b_{25}=8807546d+33705248273090$.
$\bullet$ $4341133124105a_{26}=-5385d+20420906899575$,            $33705248273090a_{26}+4341133124105b_{26}=888730d+19138455539665$.
$\bullet$ $20420906899575a_{27}=-21329d+11628413963866$,        $19138455539665a_{27}+20420906899575b_{27}=432256d+14371072675706$.
$\bullet$ $11628413963866a_{28}=-10322d+27536499676132$,        $14371072675706a_{28}+11628413963866b_{28}=2230673d+20483449818325$.
$\bullet$ $27536499676132a_{29}=-16214d+11871313871146$,        $20483449818325a_{29}+27536499676132b_{29}=10491877d+23841147446157$.
$\bullet$ $11871313871146a_{30}=1168d+36708968638606$,             $23841147446157a_{30}+11871313871146b_{30}=583202d+29824187122383$.
$\bullet$ $36708968638606a_{31}=27670d+35292993605984$,           $29824187122383a_{31}+36708968638606b_{31}=6287182d+30828612734725$.
$\bullet$ $35292993605984a_{32}=33791d+32099367907223$,           $30828612734725a_{32}+35292993605984b_{32}=2928174d+15667131309718$.
$\bullet$ $32099367907223a_{33}=31502d+8686497691888$,             $15667131309718a_{33}+32099367907223b_{33}=1722251d+13069916858845$.
$\bullet$ $8686497691888a_{34}=17842d+32178145784978$,             $13069916858845a_{34}+8686497691888b_{34}=1025368d+38021671372563$.
$\bullet$ $32178145784978a_{35}=68542d+38082283375434$,           $38021671372563a_{35}+32178145784978b_{35}=5937837d+11732682959923$.
$\bullet$ $38082283375434a_{36}=100145d+35965182833267$,         $11732682959923a_{36}+38082283375434b_{36}=6202247d+37211608563518$.
$\bullet$ $35965182833267a_{37}=106375d+33169240097520$,         $37211608563518a_{37}+35965182833267b_{37}=8936948d+26020836879171$.
$\bullet$ $33169240097520a_{38}=99259d+17708055977779$,           $26020836879171a_{38}+33169240097520b_{38}=10150666d+32544641385778$.
$\bullet$ $17708055977779a_{39}=70781d+30770548150680$,           $32544641385778a_{39}+17708055977779b_{39}=2202939d+36915064492287$.
$\bullet$ $30770548150680a_{40}=155955d+11904176271315$,         $36915064492287a_{40}+30770548150680b_{40}=2562988d+165258901717$.
$\bullet$ $11904176271315a_{41}=64624d+25508831080099$,           $165258901717a_{41}+11904176271315b_{41}=1347305d+5907415462461$.
$\bullet$ $25508831080099a_{42}=143454d+872140076252$,             $5907415462461a_{42}+25508831080099b_{42}=790654d+34225185690977$.
$\bullet$ $872140076252a_{43}=5137d+10987048845161$,                  $34225185690977a_{43}+872140076252b_{43}=339647d+26138731918415$.
$\bullet$ $10987048845161a_{44}=74345d+8152289751913$,             $26138731918415a_{44}+10987048845161b_{44}=1871482d+38377823723539$.
$\bullet$ $8152289751913a_{45}=56296d+17399000858672$,             $38377823723539a_{45}+8152289751913b_{45}=295750d+32817075202061$.
$\bullet$ $17399000858672a_{46}=140874d+13424492118010$,         $32817075202061a_{46}+17399000858672b_{46}=2189576d+33163475056782$.
$\bullet$ $13424492118010a_{47}=116689d+22068166729399$,         $33163475056782a_{47}+13424492118010b_{47}=705436d+24070914532262$.
$\bullet$ $22068166729399a_{48}=266604d+20874064459588$,         $24070914532262a_{48}+22068166729399b_{48}=292240d+15318085049644$.
$\bullet$ $20874064459588a_{49}=267012d+25167892634080$,         $15318085049644a_{49}+20874064459588b_{49}=789145d+30753673516637$.
$\bullet$ $25167892634080a_{50}=384789d+20538358542749$,         $30753673516637a_{50}+25167892634080b_{50}=1701637d+23065592907429$.
$\bullet$ $20538358542749a_{51}=591526d+21236595367747$,         $23065592907429a_{51}+20538358542749b_{51}=3306393d+31259701520158$.
$\bullet$ $21236595367747a_{52}=628169d+663951808753$,             $31259701520158a_{52}+21236595367747b_{52}=1963739d+37377022803680$.
$\bullet$ $663951808753a_{53}=20667d+11352564579816$,               $37377022803680a_{53}+663951808753b_{53}=1565314d+35652369557844$.
$\bullet$ $11352564579816a_{54}=408337d+12677902910401$,         $35652369557844a_{54}+11352564579816b_{54}=6095135d+24252482581547$.
$\bullet$ $12677902910401a_{55}=632652d+37493335194470$,         $24252482581547a_{55}+12677902910401b_{55}=2876153d+28715685383254$.
$\bullet$ $37493335194470a_{56}=1912828d+27624385992608$,       $28715685383254a_{56}+37493335194470b_{56}=11211960d+33470792628624$.
$\bullet$ $27624385992608a_{57}=1505206d+34038147456710$,       $33470792628624a_{57}+27624385992608b_{57}=6487920d$.
@@
$z\equiv$ $\left(0+\sqrt{d}\right)^{3}\left(2+\sqrt{d}\right)^{6}\left(0+\sqrt{d}\right)^{2}\left(1+\sqrt{d}\right)^{3}\left(2+\sqrt{d}\right)^{6}\left(0+\sqrt{d}\right)^{2}\left(0+\sqrt{d}\right)^{0}\left(1+\sqrt{d}\right)^{2}\left(0+\sqrt{d}\right)^{3}\left(2+\sqrt{d}\right)^{6}\left(2+\sqrt{d}\right)^{2}\left(0+\sqrt{d}\right)^{1}$
$\left(2+\sqrt{d}\right)^{5}\left(2+\sqrt{d}\right)^{7}\left(2+\sqrt{d}\right)^{1}\left(1+\sqrt{d}\right)^{3}\left(2+\sqrt{d}\right)^{2}\left(2+\sqrt{d}\right)^{3}\left(0+\sqrt{d}\right)^{1}\left(2+\sqrt{d}\right)^{3}\left(1+\sqrt{d}\right)^{7}\left(0+\sqrt{d}\right)^{3}\left(1+\sqrt{d}\right)^{2}\left(1+\sqrt{d}\right)^{4}$
$\left(1+\sqrt{d}\right)^{0}\left(0+\sqrt{d}\right)^{3}\left(1+\sqrt{d}\right)^{3}\left(0+\sqrt{d}\right)^{3}\left(0+\sqrt{d}\right)^{1}\left(0+\sqrt{d}\right)^{3}\left(1+\sqrt{d}\right)^{7}\left(0+\sqrt{d}\right)^{1}\left(1+\sqrt{d}\right)^{2}\left(1+\sqrt{d}\right)^{3}\left(1+\sqrt{d}\right)^{2}\left(0+\sqrt{d}\right)^{2}$
$\left(1+\sqrt{d}\right)^{7}\left(2+\sqrt{d}\right)^{1}\left(1+\sqrt{d}\right)^{6}\left(2+\sqrt{d}\right)^{5}\left(1+\sqrt{d}\right)^{5}\left(2+\sqrt{d}\right)^{3}\left(2+\sqrt{d}\right)^{3}\left(1+\sqrt{d}\right)^{1}\left(1+\sqrt{d}\right)^{7}\left(2+\sqrt{d}\right)^{0}\left(0+\sqrt{d}\right)^{0}\left(1+\sqrt{d}\right)^{4}$
$\left(1+\sqrt{d}\right)^{2}\left(2+\sqrt{d}\right)^{7}\left(0+\sqrt{d}\right)^{0}\left(2+\sqrt{d}\right)^{7}\left(0+\sqrt{d}\right)^{2}\left(2+\sqrt{d}\right)^{6}\left(2+\sqrt{d}\right)^{7}\left(2+\sqrt{d}\right)^{6}\left(2+\sqrt{d}\right)^{6}$
@@
It follows from Claims 2 and 4 that $\eta\in\mathcal{O}_{K}^{\times}\setminus\{1\}$. Since $\varepsilon^{-d}<\eta<\varepsilon^{d}$, there is some $k\in\mathbb{Z}\setminus\{0\}$ such that $-d<k<d$ and $\eta=\varepsilon^{k}$. Set ${\rm ord}\left(\varepsilon\mathcal{O}_{d}^{\times}\right)=\min\{m\in\mathbb{N}\mid\left(\varepsilon\mathcal{O}_{d}^{\times}\right)^{m}=\mathcal{O}_{d}^{\times}\}$. Observe that ${\rm ord}\left(\varepsilon\mathcal{O}_{d}^{\times}\right)=\left(\mathcal{O}_{K}^{\times}:\mathcal{O}_{d}^{\times}\right)\mid d$ (because $\mathcal{O}_{K}^{\times}/\mathcal{O}_{d}^{\times}$ is generated by $\varepsilon\mathcal{O}_{d}^{\times}$ and $d\in\mathbb{P}$ is ramified in $\mathcal{O}_{K}$). Since $\left(\varepsilon\mathcal{O}_{d}^{\times}\right)^{k}=\eta\mathcal{O}_{d}^{\times}=\mathcal{O}_{d}^{\times}$ by Claim 3, we have that ${\rm ord}\left(\varepsilon\mathcal{O}_{d}^{\times}\right)\mid k$, and thus $\left(\mathcal{O}_{K}^{\times}:\mathcal{O}_{d}^{\times}\right)=1$ (since $d\in\mathbb{P}$). We infer that $\varepsilon\in\mathcal{O}_{d}$, and hence $d\mid y$.∎(Theorem 1.3)